OpenStudy (anonymous):

solve the following system of equations algebraically : 5=y-x 4x^2=-17x=y=4

myininaya (myininaya):

this looks weird to me

myininaya (myininaya):

are you sure you typed it correctly?

OpenStudy (anonymous):

5=y-x 4x^2=-17x+y+4

myininaya (myininaya):

ok that looks better

myininaya (myininaya):

\[5+x=y\] is the fist equation plug this into the second equation

myininaya (myininaya):

\[4x^2=-17x+(5+x)+4\] put everything on one side you can use the quadratic formula if you want

OpenStudy (anonymous):

what do you mean by putting it all on one side

myininaya (myininaya):

\[4x^2=-16x+9\] \[4x^2+16x-9=0\] thats what i mean